OpenStudy (anonymous):

greatest common divisor of 6 and 4

OpenStudy (anonymous):

2

jabez177 (jabez177):

2.

OpenStudy (er.mohd.amir):

6=2*3 and 4=2*2 then GCD is 2.

OpenStudy (anonymous):

thanks